OpenStudy (anonymous):

Find k so that the following function is continuous on any interval. h(x)= kcosx if 0≤x≤9pi h(x)= 12-x if 9pi

OpenStudy (anonymous):

please someone help me! I've been stuck for the longest time and the answer that I get is incorrect :(

myininaya (myininaya):

@ecarreno The last inequality doesn't make sense to me 9pi<0?

myininaya (myininaya):

maybe you meant x>9pi?

OpenStudy (anonymous):

@myininaya Yes oops sorry.

myininaya (myininaya):

If so, it makes sense to me... And what you need is the following: If h is continuous at x=9pi then you have the following: a) \[\lim_{x \rightarrow 9 \pi} h(x) \text{ exists } \] For this to exist you must have the left limit equal the right limit. b) \[h(9 \pi) \text{ exists }\] and finally c) You want the number you get from a to equal the number you got from b. Let's start off with the left limit equaling the right limit

myininaya (myininaya):

\[\lim_{x \rightarrow 9\pi^+}h(x)=\lim_{x \rightarrow 9\pi^-}h(x)\]

myininaya (myininaya):

Evaluate both sides first. Then solve the equation.

OpenStudy (anonymous):

I don't understand which equation? this is how I was solving it, please tell me if it's the correct way... kcosx=12-x kcos9pi=12-9pi -k=12-9pi k=-12+9pi

myininaya (myininaya):

\lim_{x \rightarrow 9\pi^+}h(x)=\lim_{x \rightarrow 9\pi^-}h(x) I was talking about this equation.... But yeah you went about it correct you can plug in 9pi because both of the functions are continuous at 9pi

myininaya (myininaya):

\[\lim_{x \rightarrow 9\pi^+}h(x)=\lim_{x \rightarrow 9\pi^-}h(x)\] To evaluate the right limit you get 12-9pi To get the left limit you get kcos(9pi) So \[\lim_{x \rightarrow 9\pi^+}h(x)=\lim_{x \rightarrow 9\pi^-}h(x)\] we can rewrite this equation as: 12-9pi=kcos(9pi) Then solve for k

OpenStudy (anonymous):

Okay my answer is -12+9pi after I solve for k and the online calc book I am using tells me that it is incorrect.

myininaya (myininaya):

Did they write 9pi-12?

myininaya (myininaya):

Because that is the same thing

OpenStudy (anonymous):

No, it won't show me the answer unless I use up all of my attempts.. I am down to my last attempt

myininaya (myininaya):

I don't know if they want an approximation or not but your answer is fine in the exact form

OpenStudy (anonymous):

it was 9pi+12, thank you:)

myininaya (myininaya):

I think the computer or the test maker made an error in that answer

OpenStudy (anonymous):

oops I meant 9pi-12

myininaya (myininaya):

It should be 9pi-12

myininaya (myininaya):

But you know that is the exact answer you put

myininaya (myininaya):

Just in a different order

myininaya (myininaya):

9pi-12 is -12+9pi

OpenStudy (anonymous):

yeah, but i guess the online book wants it in a specific order..

myininaya (myininaya):

Ok. Coolness.